OpenStudy (deadshot):

How do I solve logv12 (x + 1) - logv12 4 = logv12 16?

OpenStudy (deadshot):

\[\log_{12} (x+1) - \log_{12} 4=\log_{12} 16\]

OpenStudy (nathan917):

Solve log12 (x + 1) – log12 4 = log12 16 log12 ((x+1)/4) = log12 16 ((x+1)/4) = 16 (x+1) = 64 x=63

OpenStudy (nathan917):

Is there any more? @DeadShot ?

OpenStudy (deadshot):

No, that's all of them, thanks for the help!

OpenStudy (nathan917):

Fan me if you think you will have more.

OpenStudy (nathan917):

*later on*

OpenStudy (deadshot):

I did, I might have more in a little while.

OpenStudy (nathan917):

Ok. np.